What is the change in internal energy of a system if the surroundings absorb 145 J of heat from the system while doing 98 J of work on the system?


1
Expert's answer
2021-11-12T01:19:05-0500

This is the first law of thermodynamics

"\Delta U=A+Q=A-Q=98-145=-47" J